OpenStudy (anonymous):

Simplify the expression by using a double angle formula and or half angle formula. Sin2degrees/1+cos2degrees

OpenStudy (anonymous):

\[ \frac{ \sin 2\alpha }{ 1+\cos 2\alpha }=\frac{ 2\sin \alpha \cos \alpha }{ 1+(2\cos^{2} \alpha -1) }=\frac{ 2\sin \alpha \cos \alpha }{ 2\cos^{2} \alpha }=\frac{ \sin \alpha }{ \cos \alpha }=\tan \alpha\]

OpenStudy (anonymous):

tan1